OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y - z = 6 3x - y - 2z = 4 3x - 2y + 3z = 13 A)x + 5y + 6 B)x + 5y - 6 C)-x - 5y + 6 D)-x - 5y - 6

OpenStudy (anonymous):

@iGreen

OpenStudy (mathstudent55):

Take the top equation and solve it for z.

OpenStudy (mathstudent55):

x + 5y - z = 6 You want z by itself. First, add 6 to both sides. Then, add z to both sides. What do you get?

OpenStudy (igreen):

@kellyconnelly I wanna see how you really do it..so I'm gonna watch @mathstudent55

OpenStudy (anonymous):

x + 5y - z = 6 +z -6 x + 5y -6 = z

OpenStudy (anonymous):

this question is alot easier than I thought.. ahah

OpenStudy (igreen):

Nevermind..Lol.

OpenStudy (anonymous):

is B correct

OpenStudy (igreen):

I guess so..

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

i wanna double check hehhe

OpenStudy (mathstudent55):

Great job. You have already done it. z = x + 5y - 6 B is the answer, and you're done.

OpenStudy (mathstudent55):

Not all math questions need to be hard. :-)